OpenStudy (anonymous):

Given 2x - y = 7, solve for y.

OpenStudy (the_fizicx99):

Subtract both sides 2x, 2x - y = 7 -2x -2x -y = -2x + 7 /-1 /-1 /-1 divide every term by negative 1 to make the y positive, y = 2x - 7